OpenStudy (anonymous):

x^6 + 26x^3 - 27 is it equal to : (x^3 + 27)(x^3 - 1) ?

hartnn (hartnn):

lets check x^6 +27x^3-x^3-27 = 0 x^3 (x^3+27) -1 (x^3+27) =0 (x^3+27)(x^3-1) = 0 correct! but both of these can be further factorized, do you know the relations ?

OpenStudy (anonymous):

How can I factorize it further? :)

hartnn (hartnn):

\(\large a^3+b^3 = (a+b)(a^2-ab+b^2) \\ \large a^3-b^3 = (a-b)(a^2+ab+b^2)\)

OpenStudy (anonymous):

Oh. Thankyou! :)

OpenStudy (anonymous):

So it will be (x+3)(x^2-3x+9) ?

hartnn (hartnn):

correct for x^3+27 what about x^3-1 ??

OpenStudy (anonymous):

I'm still trying to figure it out :))

hartnn (hartnn):

\( \\ \large a^3-b^3 = (a-b)(a^2+ab+b^2) \\a=x, b=1 \\ \large x^3-1^3=...?\)

OpenStudy (anonymous):

will it be (x-1)(x^2-x+1) ?

hartnn (hartnn):

in the term x^2-x+1 will it be -x or +x ? take a look at the formula...

OpenStudy (anonymous):

Oh. yeah XD sorry. My mistake. Thankyouuuuu!!!

hartnn (hartnn):

welcome ^_^